OpenStudy (anonymous):

A little car has a maximum acceleration of 2.57 m/s 2. What is the maximum acceleration of the little car if it tows another car that has the same mass?

OpenStudy (anonymous):

(ma)/m = a a = a (ma)/2m = a a/2 = a 2.57 / 2 = 1.285 m/s^2

OpenStudy (anonymous):

I wrote that in a strange way so if you don't understand I can clean it up if needed.